OpenStudy (anonymous):

\[1+\cos \alpha+\cos ^{2}\alpha+............=2-\sqrt{2}\] find the value of alpha.

OpenStudy (anonymous):

isn't this an infinite geometric series with r = cosx?

OpenStudy (anonymous):

here are all the solutions https://www.wolframalpha.com/input/?i=2+-+sqrt%282%29+%3D+1+%2F+%281+-+cos%28x%29%29

mathslover (mathslover):

\(\sf{1 + \cos \alpha + \cos^2\alpha + \cos^3 \alpha + ..... = 2-\sqrt{2} \\ \cos \alpha + \cos^2 \alpha +\cos^3 \alpha + .... = 1 - \sqrt{2} \\ \cos \alpha ( 1 + \cos \alpha + \cos^2\alpha+ ....) = 1 - \sqrt{2} \\ \cos \alpha (2 - \sqrt{2} ) = 1-\sqrt{2} \\ \cos \alpha = \cfrac{1-\sqrt{2}}{2-\sqrt{2}}\\ \cos \alpha = \cfrac{(1-\sqrt{2})(2+\sqrt{2})}{2} \\ \cos \alpha = \cfrac{2 + \sqrt{2} -2\sqrt{2} - 2}{2} \\ \cos \alpha = \cfrac{-\sqrt{2}}{2} \\ }\)

OpenStudy (mathmale):

@sourwing is right in suggesting that this is a geometric series with common ratio \[\cos \alpha\]. Looking at the given series as it now stands, I see that the initial value, a, is just 1. The sum of a geom. series with intial value a and common ratio r is \[S=\frac{ a }{ 1-r }\] and we are told that the sum of this particular series is \[2-\sqrt{2}\]

mathslover (mathslover):

Yes mathmale. But, I am confused where I went wrong?

OpenStudy (mathmale):

thus, you may equate S = a / (1-r) to \[2-\sqrt{2}\] With a =1 and r = cos alpha, \[S=\frac{ 1 }{ 1-\cos \alpha }=2-\sqrt{2}\]

OpenStudy (mathmale):

Think you can solve this for cos alpha? for alpha alone?

OpenStudy (mathmale):

Note that alpha must NOT be 0, since cos alpha would then be 1, and we'd have division by zero.

OpenStudy (mathmale):

I need to hit the sack pretty pronto! Hint: I've obtained two solutions, and one of them is 3Pi/4. By all means check by substituting 3Pi/4 for alpha in the original equation for the sum of the series. Is the sum actually 2-\sqrt{2}? Great working with you. If you need more input from me, just private message me in the morning and I'll try to help you further. Good night.

OpenStudy (mathmale):

By all means please try solving for alpha yourself and identifying the other root (besides 3Pi/4).

mathslover (mathslover):

Okay, thanks a lot Math! :) Good Night